(10 pts) The iron content of a well water sample is determined by UV-Vis spectroscopy. A 10.00 mL aliquot of well water is transferred to a 100.0-mL volumetric flask and 20 mL of 0.10 M nitric acid is added to oxidize iron (II) to iron (II). Next 30 mL of a 0.1 M sodium thiocyanate solution is added to form the deep red complex iron (II) thiocyanate. The solution is diluted to volume with DI water. A reagent blank is similarly prepared. If the absorbance of the iron thiocyanate solution is 0.5483 and the reagent blank absorbance is 0.0282, calculate the iron concentration for the water sample. The molar absorptivity for iron (III) thiocyanate is 3200 M cm1. Assume a path length of 1 cm and complete reaction with the thiocyanate. Fe (aq) + SCN. (aq) ㈠ FeSCN2+ (aq)
Answer #1

absorbance of solution = 0.5483

absorbance of blank = 0.0282

Absorbance of Fe(SCN)2+ solution = 0.5483 - 0.0282 = 0.5201

molar absorptivity = 3220 M-1.cm-1

path length = 1 cm

concentration of Fe(SCN)2+ in dilute solution = 0.5201/3220 M-1.cm-1 x 1 cm

                                                                         = 1.615 x 10^-4 M

molarity of Fe in original 10 ml water sample = 1.615 x 10^-4 M x 100/10 = 1.615 x 10^-3 M